what is the frequency of an acrobat swinging on a trapeze if the cables are approximately 2.5 m long?
1
Expert's answer
2021-10-04T13:34:06-0400

The frequency is


"f=\frac{1}{T}=\frac{1}{2π}\sqrt{\frac{g}{l}}=0.32\text{ Hz}"